OpenStudy (anonymous):

isnt this D? Which function below is the inverse of f(x) = x2 − 36? x squared over 36 ±6square root of x 1 over quantity x squared minus 36 ±square root of the quantity x plus 36

OpenStudy (zale101):

\(y=x^2-36\) switch the x and y

OpenStudy (zale101):

@hannah120899 you there?

OpenStudy (anonymous):

oh okay and yes! So its not D

OpenStudy (zale101):

Yes, it's D.

OpenStudy (anonymous):

oh okay i thought so lol thanks

OpenStudy (zale101):

\(y=x^2-36\) Switch the x and y \(x=y^2-36\) Add 36 to both sides \(x+36=y^2\) Square root to both sides \(\pm \sqrt{x+36}=y\)